The current price of a non-dividend-paying stock is $160. Over the next year it is expected to rise to $176 or fall to $154. Assume the risk free rate is 5% per year. An investor buys a European call option with a strike price of $162 per share. Assume that the option is written on 100 shares of stock. What stock position should the investor take today so that she would hold a riskless portfolio if it was combined with the long call option position?
Answer #1

Delta of Call=(MAX(176-162,0)-MAX(154-162,0))/(176-154)=0.63636364

Short Delta*100=0.63636364*100=64 shares